six-month Treasury bill and a nine-month bill both sell at a discount of 8%.

a-1. Calculate the annual yield of the six-month Treasury bill. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) Annual yield %

a-2. Calculate the annual yield of the nine-month Treasury bill. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) Annual yield %

Answers are NOT 8.16 or 8.08 respectively.
Answer #1

A-1

Assume par Value is $1000

Discount rate is 8%

Number of months = 6

In years (n)= 6/12 = 0.5

Price of T-bill would be = Face Value *(1-(Discount Rate*n))

=1000*(1-(8%*0.5))

=$960

Annual Yield of T-bill formula = (( face Value/price of T-bill)^(1/n))-1

=((1000/960)^(1/0.5))-1

=0.08506944444 or 8.51%

Annual yield of 6 month T-bill is 8.51%

A-2

Assume par Value is $1000

Discount rate is 8%

Number of months = 9

In years (n)= 9/12 = 0.75

Price of T-bill would be = Face Value *(1-(Discount Rate*n))

=1000*(1-(8%*0.75))

=$940

Annual Yield of T-bill formula = (( face Value/price of T-bill)^(1/n))-1

=((1000/940)^(1/0.75))-1

=0.08599925802 or 8.60%

So Annual Yield of 9 month T-bill is 8.60%